Find the general solution of these differential equations.

7) ydy/dx =1/ √(x) 8) dy/dx = 4x√(1-y2) 9) dy/dx = x(4+y2)




1
Expert's answer
2021-09-16T00:46:51-0400

Let us find the general solution of these differential equations.


7) Let us solve the equation "y\frac{dy}{dx} =\frac{1}{\sqrt{x}}" which is equivalent to "ydy =\frac{dx}{\sqrt{x}}." It follows that "\int ydy =\int\frac{dx}{\sqrt{x}}," and we conclude that the general solution is "\frac{y^2}2=2\sqrt{x}+C."


8) Let us solve the equation "\frac{dy}{dx} = 4x\sqrt{1-y^2}" which is equivalent to "\frac{dy}{\sqrt{1-y^2}} = 4xdx." It follows that "\int\frac{dy}{\sqrt{1-y^2}} = 4\int xdx," and therefore, the general solution is "\arcsin y=2x^2+C."


9) Let us solve the equation "\frac{dy}{dx} = x(4+y^2)" which is equivalent to "\frac{dy}{4+y^2} = xdx." It follows that "\int\frac{dy}{4+y^2} = \int xdx," and we conclude that the general solution is "\frac{1}2\arctan \frac{y}2=\frac{x^2}2+C."